OpenStudy (anonymous):

New Problem: Need Help Bad A 100 lbs of radioactive substance decays to 35 lbs in 10 years. How much remains in 27 years? I think if i could be taught how to find the value for Half-Life, that i could solve this problem. Please help someone.

OpenStudy (anonymous):

@phi if you are free to help i'd appreciate it. @Titanic12 if you dont mind i really need help with this one part of this problem.

OpenStudy (phi):

you could use the equaton \[ A_t= A_0 \left(\frac{1}{2} \right)^{\frac{t}{\lambda}} \] where A0 is the starting amount, t is the number of years, and lambda is the half-life

OpenStudy (anonymous):

so its \[35=100(1/2)^{10/h}\]

OpenStudy (phi):

they did not give the half-life, so we would have to figure it out, using the info they gave. we could just use \[ A_t = A_0 e^{kt} \] where k is an unknown constant. We would have to solve for k using the info...

OpenStudy (phi):

yes. can you find h ?

OpenStudy (anonymous):

and the h being the half life, thats the part i cant get. im not sure how to find half life

OpenStudy (anonymous):

cause i had gotten the same equation, this is the part i got stuck on

OpenStudy (phi):

first divide both sides by 100 .35 = (1/2)^(10/h)

OpenStudy (phi):

now take the log of both sides write log "in front of" each side.

OpenStudy (anonymous):

ok i got that if im doing it right so far im to the point where this is the equation i have: \[\log(0.35)=\frac{ -10 }{ h }\log(2)\]

OpenStudy (phi):

ok

OpenStudy (anonymous):

then i believe next step is to divide both sides by log(2) so \[\frac{ \log(0.35) }{ \log(2)}=\frac{ -10 }{ h }\]

OpenStudy (anonymous):

im not sure how to get the h by itself now

OpenStudy (phi):

you could "flip" both sides

OpenStudy (phi):

and then multiply both sides by -10

OpenStudy (anonymous):

so it would be \[\frac{ \log2 }{ \log0.35 }=\frac{ h }{ -10 }\] then \[-10[\frac{ \log(2) }{ \log(0.35) }]=h\]

OpenStudy (anonymous):

and i just put that into my calculator to get a number?

OpenStudy (phi):

yes and it would be more convenient to find 1/h = k so we can use \[ A_t= A_0\left( \frac{1}{2}\right)^{kt} \]

OpenStudy (anonymous):

ok so i put it in my calculator and i get h=6.602520221 or h=6.6

OpenStudy (anonymous):

what do you mean it would be more convenient to find that? like i should use that equation instead of what i just did?

OpenStudy (phi):

with exponents, it's always good to keep lots of digits. when I suggest using k= 1/h that is because it's (slightly) simpler to do t*k rather than t/h.

OpenStudy (anonymous):

ooh alright i see

OpenStudy (anonymous):

so is that equation A = A(1/2)^kt the same as At=A0 e^kt? just wrote it quickly without formatting it

OpenStudy (phi):

related but different if we used the second form (which does not show the half-life), the math would have been 0.35 = e^10k ln(0.35) = 10 k k= ln(0.35)/10 it's a bit faster to find the exponent.

OpenStudy (anonymous):

oh alright let me write that out and see if i understand it

OpenStudy (phi):

the "k" s will have different values for the two forms. In the first 1/k is the half-life in the second, the half-life would be ln(0.5)/k where k is ln(0.35)/10

OpenStudy (anonymous):

ok yeah so i got h=6.6 again

OpenStudy (phi):

can you finish the problem?

OpenStudy (anonymous):

yes but just real quick, i did \[\frac{ \ln(0.5) }{ \frac{ \ln(0.35 }{ 10 } }\] but where is the ln(0.5) from?

OpenStudy (anonymous):

sorry for all the quesitons, just trying to understand whats going on

OpenStudy (phi):

all exponentials are related. we use this property: \[ e^{\ln(x)} = x\] to write 1/2 as \[ 0.5 = e^{\ln(0.5)} \] and the original equation \[ A_t= A_0 (0.5)^\frac{t}{h} \\ = A_0 \left( e^{\ln(0.5)}\right)^\frac{t}{h} \\ = A_0e^{\frac{\ln(0.5)}{h} \cdot t} \]

OpenStudy (anonymous):

oooh ok i see where that comes from now! thank you so much!!

OpenStudy (anonymous):

i got my final answer of how many lbs after 27 years remains is 5.87471 lbs

OpenStudy (anonymous):

pretty sure i did that right.

OpenStudy (phi):

looks good

OpenStudy (anonymous):

thank you so much phi, your a huge help!!